# Evaluating Large Language Models as Generative User Simulators for Conversational Recommendation

### v2

@@ 1 Introduction
(T1) ItemsTalk DI, IH ReDial, Reddit, IMDB
A person mentions Concussion (2015) and Jerry Maguire (1996) in a conversation about movies and proceeds to mention 2 more. What would these 2 movies be?
(T2) BinPref DI, DI + PP MovieLens
Pretend to be Ms. Guzman. You watched the movie Whiplash (2014). Did you like the movie? Answer Yes or No. Don’t say anything else.
(T3) OpenPref DI, DI + PP IMDB
Pretend to be Mr. Li. You watched the movie The Bellboy (1960). What are your thoughts on this movie? Answer should not exceed 809 characters.
(T4) RecRequest Vanilla LLM Reddit
Generate a movie recommendation request. Include the following movies in your text: Oldboy (2003), Memento (2000). Length of the request is approximately 374 characters.
(T5) Feedback Vanilla LLM Reddit
In the following conversation … If the recommendation is coherent to your request, answer Accept. If the recommendation is incoherent to your request, answer Reject.
@@ 3 Methods
BinPref prompts the simulator to act as if one has interacted with an item, and asks whether one has positive opinions on it. We sample two groups of $200$ movies from MovieLens: frequent ($\geq 5000$ ratings) and infrequent ($\leq 500$ ratings). This is to observe if the simulators reflect human preferences better on frequent movies. We additionally sample $300$ movies regardless of frequency. The distribution of average rating ranges from 1 to 5. For each movie, we run $100$ simulators to output a binary preference and get the proportion of ‘Yes’ answers (i.e., positive rate). We compute the Pearson correlation coefficient between the average rating and positive rate.
@@ 4 Experiments
Movies showing extreme loneliness or depression. I have watched Taxi Driver (1976) and Joker (2019) and would like to see more similar movies showing loneliness or depression.
Looking for a gripping psychological thriller similar to Taxi Driver (1976) or Joker (2019)? Seeking a movie that delves into the mind of complex characters?
Movies about conspiracies, lies, and finding the truth like Memento (2000) and Fight Club (1999)?. Espcially ones that have big plot twists.
Looking for mind-bending thrillers like Fight Club (1999) and Memento (2000). Any suggestions? Need gripping plots that leave me questioning reality!
@@
Human: ‘Movies about alcoholism’, ‘Space movies?’, ‘Movies about redemption’, ‘Inspirational movies’, ‘Good biography movies’, ‘Impactful endings?’, ‘Growth mindset versus Fixed Mindset’, ‘Rock climbing movies’, ‘Movies about nihilism’ gpt-3.5-turbo: ‘Movie recommendation?’, ‘Need movie suggestions’, ‘Need movie recs!!’, ‘Need movie recommendations’, ‘Movie recs?’, ‘Movie recommendations?’ gpt-4: ‘Got recs?’ text-davinci-003: ‘Recommend a movie’, ‘Cheerful movies?’, ‘Recommend me!’
@@
Request: I’m in a state of life rn that i really want/need movies with the Main Character being a loner or alone in general, he doesn’t have to be alone the entire movie, but just portrayal of a good loner main character. Recommendation: Nightcrawler (2014) Feedback from simulator: Reject.
@@ A.1 Dataset statistics
The MovieLens dataset contains ratings of $62,000$ movies rated by $162,000$ users. We sample $200$ movies rated by more than $5000$ users (frequent movies) and $200$ movies rated by less than $500$ and more than $50$ users (infrequent movies). We also sample $300$ movies regardless of rating frequency.
@@ A.2.5 Feedback
Figure 12: Preference trends of humans and simulators, on movies sampled across all frequency levels. We did not obtain results for text-davinci-003, since by the time we ran these experiments, the model was deprecated.
Table 8: Correlation coefficient between human and simulator preferences, where items are sampled across all frequency levels.
@@ A.3.2 Results from BinPref
We compare the trends of average rating and positive rate (proportion of simulators that answered that they liked the movie). Results for items across all frequency levels are shown in Table 8 and Figure 12. Plots for frequent items are shown in Figure 13 and infrequent items in Figure 14.

### v4

@@ 1 Introduction
(T1) ItemsTalk DI, IH ReDial, Reddit, IMDB A person mentions Concussion (2015) and Jerry Maguire (1996) in a conversation about movies and proceeds to mention 2 more. What would these 2 movies be?
(T2) BinPref DI, DI + PP MovieLens Pretend to be Ms. Guzman. You watched the movie Whiplash (2014). Did you like the movie? Answer Yes or No. Don’t say anything else.
(T3) OpenPref DI, DI + PP IMDB Pretend to be Mr. Li. You watched the movie The Bellboy (1960). What are your thoughts on this movie? Answer should not exceed 809 characters.
(T4) RecRequest Vanilla LLM Reddit Generate a movie recommendation request. Include the following movies in your text: Oldboy (2003), Memento (2000). Length of the request is approximately 374 characters.
(T5) Feedback Vanilla LLM Reddit In the following conversation … If the recommendation is coherent to your request, answer Accept. If the recommendation is incoherent to your request, answer Reject.
@@ 3 Methods
BinPref prompts the simulator to act as if one has interacted with an item, and asks whether one has positive opinions on it. We sample two groups of $200$ movies from MovieLens: frequent ($\geq 5000$ ratings) and infrequent ($\leq 500$ ratings). This is to observe if the simulators reflect human preferences better on frequent movies. The distribution of average rating ranges from 1 to 5. For each movie, we run $100$ simulators to output a binary preference and get the proportion of ‘Yes’ answers (i.e., positive rate). We compute the Pearson correlation coefficient between the average rating and positive rate.
@@ 4 Experiments
Movies showing extreme loneliness or depression. I have watched Taxi Driver (1976) and Joker (2019) and would like to see more similar movies showing loneliness or depression. Looking for a gripping psychological thriller similar to Taxi Driver (1976) or Joker (2019)? Seeking a movie that delves into the mind of complex characters?
Movies about conspiracies, lies, and finding the truth like Memento (2000) and Fight Club (1999)?. Espcially ones that have big plot twists. Looking for mind-bending thrillers like Fight Club (1999) and Memento (2000). Any suggestions? Need gripping plots that leave me questioning reality!
@@
{mdframed}
  • Human: ‘Movies about alcoholism’, ‘Space movies?’, ‘Movies about redemption’, ‘Inspirational movies’, ‘Good biography movies’, ‘Impactful endings?’, ‘Growth mindset versus Fixed Mindset’, ‘Rock climbing movies’, ‘Movies about nihilism’
  • gpt-3.5-turbo: ‘Movie recommendation?’, ‘Need movie suggestions’, ‘Need movie recs!!’, ‘Need movie recommendations’, ‘Movie recs?’, ‘Movie recommendations?’
  • gpt-4: ‘Got recs?’
  • text-davinci-003: ‘Recommend a movie’, ‘Cheerful movies?’, ‘Recommend me!’
@@
{mdframed}
  • Request: I’m in a state of life rn that i really want/need movies with the Main Character being a loner or alone in general, he doesn’t have to be alone the entire movie, but just portrayal of a good loner main character.
  • Recommendation: Nightcrawler (2014)
  • Feedback from simulator: Reject.
@@ A.1 Dataset statistics
The MovieLens dataset contains ratings of $62,000$ movies rated by $162,000$ users. We sample $200$ movies rated by more than $5000$ users (frequent movies) and $200$ movies rated by less than $500$ and more than $50$ users (infrequent movies). While we also sample $300$ movies without considering rating frequency, we observe that due to the long-tail distribution of frequencies, this random sampling result in movies with significantly lower frequencies (median: 5, mode: 1 appears 49/300 times). We nonetheless show the results.
@@ A.2.5 Feedback
Figure 12: Preference trends of humans and simulators. See Table 8 for correlation coefficients.
Table 8: Correlation coefficient between human and simulator preferences. Movies are randomly sampled across all frequency levels, resulting in a sample with movies with very low frequencies.
@@ A.3.2 Results from BinPref
We compare the trends of average rating and positive rate (proportion of simulators that answered that they liked the movie). Results for items randomly sampled regardless of frequency levels are shown in Table 8 and Figure 12.888These experiments were conducted after we received our paper reviews. OpenAI’s configurations might have been updated or modified in the meantime. We did not obtain results for text-davinci-003, since by the time we ran these experiments, the model was deprecated. However, these movies often have very low frequency values (median: 5, mode: 1 appears 49/300 times); the average ratings of these movies may not reflect true user preferences. Plots for frequent items are shown in Figure 13 and infrequent items in Figure 14.